Get Help

Online: Request Help
Phone: (707) 826-4357 • Hours
Email: help@humboldt [dot] edu
Walk-In: Library 101 • Hours
Reset HSU Password
System Status

Services For

System Status Archive

System Issue start Who is affected What happened
Off-Campus Internet Connection Mar 12 2016 - 06:00 Anyone on campus accessing off campus resources Firewalls are being replaced as part of a CSU-wide project **Update 1** As of 07:00 the new firewalls were up and stable. Validation testing is ongoing, but all services appear to be working both inbound and outbound
Campus Phone System Mar 12 2016 - 07:00 Hopefully no one will be affected regarding the Campus Phone & Voicemail Systems TNS needs to replace a defective inverter in Phone and Voicemail system. This is not expected to cause any problems.
Network Folders Mar 7 2016 - 14:04 People who use Folders
Other Mar 5 2016 - 13:44 Some people trying to access the myHumboldt portal ITS is investigating. [Update 15:40] One web server was identified and restarted, full functionality has been restored
Other Feb 29 2016 - 08:00 Anyone who uses the 25Live (R25) Room Reservation Software The 25Live software used for University room reservations is being upgraded to a new look and, effective Friday March 4, will be accessible off-campus from anywhere you have Internet service. When the service is back online a notice will be submitted to campus announcements.
Moodle Mar 1 2016 - 10:19 Users of Moodle Moodle has been restored.
KBOX Ticketing System Feb 25 2016 - 22:00 Anyone attempting to log into the KBOX ticketing system to view or create tickets. KBOX is experiencing delays in sending email notifications about changes to tickets. We need to reboot the system in order to correct this problem. We apologize for any inconvenience this may cause.UPDATE: KBox has been rebooted and is available.
Other Feb 25 2016 - 19:00 Users of the Bookorders website *UPDATE* The site is back up and ready for use. Moving the site to a supported version of the Drupal CMS. The same theme and page layout is being used on the new site and the move will be transparent to customers.
Other Feb 25 2016 - 17:00 Users of the Bookorders website Moving the site to a supported version of the Drupal CMS. The same theme and page layout is being used on the new site and the move will be transparent to customers.
Other Feb 25 2016 - 05:00 Users of SSO web services including OBI, CSYou, Zimride, SCLogic, Infoready, Imleagues, 25Live. Restarting shibboleth idp service. Service available.
Computer Labs Feb 23 2016 - 17:36 Slow log in times were traced to a server not responding. The server is being looked at. Server is now available. Y: drive is now available and working correctly.
Other Feb 24 2016 - 14:45 Users of SSO web services including OBI, CSYou, Zimride, SCLogic, Infoready, Imleagues, 25Live. Service out of memory. Memory increased.
Other Feb 23 2016 - 18:56 users who access the following shares on hsu-qnap01: cartlab cnrs-software cnrs-sw gis-data gisdata its-services its-software marcom sasrepo The QNAP hardware has failed, however the data is backed up to different hardware. We are not able to create a support ticket tonight but will do so first thing in the morning while traversing another path to get alternative hardware to serve the backed up data. The following shared folders are effected. cartlab cnrs-software cnrs-sw gis-data gisdata its-services its-software marcom Public sasrepo Web Please help us inform staff and faculty of the outage and assure them we will get the data served as soon as possible.
Campus Phone System Feb 23 2016 - 16:30 Campus Telephone users - message waiting indicator ** Update 1 - A reboot to the voicemail server resolved the Message Waiting Indicator light issue. Users may find the light on today notifying messages from yesterday since the indicator was not functioning for a time yesterday. ** TNS is investigating. Check back for updates.
Computer Labs Feb 23 2016 - 07:47 A server was unavailable - temporary adjustments have been made to allow Linux log in. Unknown at this time. Look for an update at 9:00 2/23
Other Feb 23 2016 - 05:00 Users of SSO web services including OBI, CSYou, Zimride, SCLogic, Infoready, Imleagues. Service restarted and available.
Other Feb 21 2016 - 05:30 All users Monthly maintenance window. UPDATE: Servers have been patched and rebooted.
CMS/Peoplesoft Feb 17 2016 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window, the following item will be installed: fix 8 to the Orientation mod (HSU_AD279_Student_Orientation_Mod).
Moodle Feb 16 2016 - 09:48 Moode Faculty and Students Due to issues with the myHumboldt Portal, Moodle also experienced intermittent access issues. Moodle service has now been restored.
Other Feb 15 2016 - 14:16 myHumboldt Portal users We are investigating. Failover to backup load balancer. UPDATE: ITS is investigating and working on the issue and will update again at 10:00.
Other Feb 11 2016 - 04:00 Students, Faculty, Staff who use OBI, CSYou, SCLogic, Imleagues, Infoready, Skyfactor, Zimride. Upgrading Shibboleth. Upgrade completed.
CMS/Peoplesoft Feb 10 2016 - 16:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window the following items will be installed: MP 19.03, an update to the Student Group Test Scores process mod (HUM_SR285_Test_Scores) and fix 2 to the AP External Ed Data Insert mod (HUM_AD413 EXTERNAL_ED_DATA_INSERT_FIX2).
Campus Data Network Feb 4 2016 - 16:44 Campus Data Network Users Link failure between CENIC Oakland aggregate switch and CENIC Humboldt CSU switch. No ETR is available.***UPDATE*** AT&T has repaired the circuit. Campus diversity to the internet restored.
Campus Data Network Feb 4 2016 - 17:40 Telonicher House users Upgrading house bandwidth connection from 15mb to 65mb
Moodle Feb 4 2016 - 17:30 Students and faculty accessing videos that have links posted in Moodle pointing to the Stream server video library We need to set up automatic backups of the Stream server as we recognize the videos stored there are a valuable resource for faculty and students.UPDATE: The stream server is available.
Other Feb 3 2016 - 13:12 Some users. We are investigating the issue. UPDATE: One of the portal web servers was not responding to requests properly. Web service on that server was restarted, and appears to have resolved the issue.
CMS/Peoplesoft Feb 3 2016 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window the following item will be installed: a small retrofit to MP 19.00.
Pharos Pay-for-Print Jan 28 2016 - 06:00 Anyone wishing to print on Pharos printers We are configuring Pharos for offline billing. Update: Pharos is back online and working normally.
Other Jan 28 2016 - 05:00 Users of web services including OBI, CSYou, Zimride, SCLogic, Imleagues. Restarting service after configuration change. Update - Service restart completed.
Network Folders Jan 26 2016 - 22:00 Users accessing their U: drive or group share Server reboot. Update- Folders has been rebooted and is now available.
feedback